# Embedding Explainable AI in NHS Clinical Safety: The Explainability-Enabled Clinical Safety Framework(ECSF)

### v1

@@ 2.1 Framework Development Approach
This study adopted a conceptual modeling approach and cross-regulatory synthesis to integrate explainability principles within existing NHS clinical-safety standards. Rather than proposing a new assurance framework, the analysis sought to align established safety, ethical, and regulatory instruments, including DCB0129 and DCB0160, the Good Machine Learning Practice (GMLP) principles, the EU Artificial Intelligence Act, and the NHS AI Assurance and T.E.S.T. frameworks [3, 4, 9, 10, 7, 8]. The objective was to identify practical points of convergence where explainability could enhance clinical risk management and governance.
The development of the framework followed three sequential stages. The first stage, regulatory mapping, involved analysing each clause of the DCB0129 and DCB0160 lifecycle in relation to corresponding principles in the GMLP, the EU AI Act, and the NHS AI Assurance frameworks. This mapping process revealed conceptual overlaps such as traceability, validation, and human oversight, as well as areas of divergence including interpretability requirements and post-market transparency obligations. The findings from this exercise provided the structural foundation for the Explainability-Enabled Clinical Safety Framework (ECSF).
@@
The third stage, framework construction, involved synthesising insights from the regulatory and technical mappings into the Explainability-Enabled Clinical Safety Framework. The ECSF introduces five explainability checkpoints aligned with the DCB lifecycle: hazard identification, validation and verification, clinical evaluation, risk control, and post-market surveillance. Each checkpoint specifies acceptable forms of explainability evidence, the corresponding DCB artefact, and the assurance purpose such as validation, traceability, or drift monitoring. Collectively, these checkpoints establish explainability as a structured, repeatable, and measurable component of clinical-safety assurance.
@@ 2.2 Analytical Perspectives
To ensure coherence between regulatory intent, technical evidence, and clinical usability, the analysis was organised around three complementary perspectives that together informed the construction of the Explainability-Enabled Clinical Safety Framework (ECSF).
The first perspective, the regulatory analysis, examined how explainability principles align with external obligations defined by the Good Machine Learning Practice (GMLP) guidance, the EU Artificial Intelligence Act, and the NHS AI Assurance frameworks [9, 10, 7]. Clauses within DCB0129 and DCB0160 were cross-referenced with these instruments to identify where explainability outputs could provide evidence of compliance with transparency, documentation, and human oversight requirements. Article 13 of the EU AI Act, which mandates transparency and provision of information to deployers, overlaps most strongly with GMLP Principles 1, 9, and 10, and to a lesser extent Principle 3. Together, these principles emphasise multidisciplinary expertise, data representativeness, clear user information, and lifecycle monitoring. This alignment ensures that explainability contributes directly to the evidentiary trail required for regulatory assurance, rather than functioning as a separate technical exercise.
@@ 2.3 Data Sources
The conceptual analysis relied on a combination of regulatory, policy, and academic sources to ensure that the Explainability-Enabled Clinical Safety Framework (ECSF) was grounded in both policy compliance and scientific evidence. The regulatory foundation was established through the review of core NHS and governmental standards, including the NHS Digital clinical-safety standards DCB0129 and DCB0160 [3, 4], the Good Machine Learning Practice (GMLP) guiding principles [9], the EU Artificial Intelligence Act [10], and the NHS AI Assurance and T.E.S.T. frameworks published between 2023 and 2024 [7, 8]. These documents provided the structural and procedural requirements against which explainability principles were mapped.
@@ 3.1 Overview
The Explainability-Enabled Clinical Safety Framework (ECSF) integrates established explainable artificial intelligence (XAI) techniques into the DCB0129 and DCB0160 lifecycle to enhance transparency, traceability, and accountability. The framework aligns the deterministic assurance logic of existing NHS safety standards with the probabilistic behaviour of AI systems and reframes explainability from a research objective into a source of verifiable clinical-safety evidence [11, 12, 13].
@@
This approach preserves compliance with existing DCB standards while operationalising the transparency and human-oversight principles articulated in the Good Machine Learning Practice (GMLP) guidance and the transparency obligations of the EU Artificial Intelligence Act [9, 10].
@@ 3.3 Explainability for Large Language Models and Generative AI
Global LLM explainability. System-level traceability can be achieved through structured documentation and lifecycle transparency. Model-card reporting summarises the model’s intended use, limitations, and performance characteristics, providing a foundation for accountability and auditability [24]. Training-data lineage documentation ensures provenance and representativeness of datasets, aligning with transparency and oversight principles in the Good Machine Learning Practice (GMLP) guidance and the EU Artificial Intelligence Act [9, 10]. Logging of fine-tuning parameters and model versioning during updates supports reproducibility and traceability throughout iterative development cycles [20]. Finally, specification of guardrail logic within safety documentation maintains human oversight and behavioural control, consistent with NHS AI Assurance and T.E.S.T. frameworks [7, 8].
@@ 3.4 Integration into Assurance Activities and Artefacts
Within the Explainability-Enabled Clinical Safety Framework (ECSF), explainability evidence is systematically integrated into the five checkpoints of the DCB lifecycle. Each checkpoint aligns with existing assurance artefacts, such as the Hazard Log, Clinical Safety Case, Clinical Evaluation Report, and Post-Market Surveillance documentation, ensuring that transparency, interpretability, and oversight are evidenced throughout the system’s lifecycle. This integration preserves DCB conformance while extending the assurance process to address the probabilistic behaviour of AI systems.
@@
Explainability-Enhanced Validation – Case-Level Explainability (Local). During model validation, local explainability techniques demonstrate whether predictions align with expected clinical reasoning. Methods such as LIME, SHAP, Integrated Gradients, and saliency mapping clarify how specific input variables contribute to individual predictions, providing case-level interpretability that complements quantitative validation metrics [17]. In large-language models (LLMs), token-level attention visualisation and rationale-trace comparison provide similar evidence. These outputs substantiate the verification and validation activities required under DCB0129 Section 3.5 and are incorporated into the Clinical Safety Case as supporting validation evidence.
Clinical Evaluation – Usability and Human Oversight. Explainability also supports human-factors assurance by enabling clinicians to interrogate model outputs during usability testing. Interactive dashboards, contrastive and counterfactual examples, and rationale summaries allow end users to review AI-generated recommendations, confirming that model reasoning aligns with clinical context and professional judgement [14, 15]. This stage aligns with Good Machine Learning Practice (GMLP) Principle 7 on the performance of the human–AI team and Principle 9 on transparency and user information. [9]. Findings are summarised within the Clinical Evaluation Report and, where interpretive or cognitive risks are identified, recorded as updates to the Hazard Log. These activities strengthen the feedback loop between technical validation and clinical usability, reinforcing interpretability as a core dimension of safety.
@@ 3.5 Illustrative Application Scenario
To demonstrate the practical application of the Explainability-Enabled Clinical Safety Framework (ECSF), a simulated example of an AI early-warning system for sepsis prediction was developed. The model analysed routinely collected vital signs and laboratory data to estimate sepsis risk. This example is illustrative and does not represent a live deployment; rather, it shows how explainability evidence can be systematically embedded within DCB0129 and DCB0160 artefacts across the assurance lifecycle.
Hazard Identification. An agggregate SHAP analysis revealed disproportionate weighting of age and white blood cell count, with under-representation of lactate and respiratory rate, indicating potential model bias and an increased risk of missed atypical presentations. This issue was recorded as a new hazard in the Hazard Log and supported by a Failure Modes and Effects Analysis (FMEA) detailing the failure mode, potential clinical impact, and proposed mitigation measures.
@@
Post-Market Surveillance. Within the Post-Market Surveillance (PMS) Plan, periodic audits of explanation stability and feature-importance drift were implemented to detect changes in model reasoning over time, corresponding to ECSF Checkpoint C5 (Post-Market Surveillance).
@@ 3.6 ECSF Alignment with Regulatory Principles and Key Outcomes
The Explainability-Enabled Clinical Safety Framework (ECSF) aligns with the primary regulatory and ethical instruments governing the safe use of artificial intelligence in the NHS. It operationalises the principles of transparency, traceability, and human oversight established in the Good Machine Learning Practice (GMLP) guidelines, the EU AI Act, and the NHS AI Assurance and T.E.S.T. frameworks, while remaining consistent with the DCB0129 and DCB0160 clinical risk management standards [9, 10, 7, 8].
@@ 4.1 Reframing Explainability as a Safety Enabler
The Explainability-Enabled Clinical Safety Framework (ECSF) provides a pragmatic mechanism for translating abstract interpretability research into operational clinical-safety practice. Traditional NHS assurance processes, developed for deterministic software, assume consistent performance under fixed parameters. However, the adaptive and probabilistic behaviour of AI systems challenges these assumptions. Even under controlled conditions, model outputs can vary across runs or datasets due to stochastic optimisation processes and sampling effects [31, 2]. This inherent variability complicates verification and reproducibility, requiring new assurance mechanisms that make uncertainty observable and manageable rather than eliminated.
Explainability provides this mechanism by transforming AI from an opaque computational process into an auditable system. Rather than seeking to remove uncertainty, the ECSF manages it through structured visibility. Explainability checkpoints embedded within the DCB0129 and DCB0160 lifecycle document, interpret, and contextualise model behaviour, converting opacity into evidence. This allows Clinical Safety Officers (CSOs) to trace model reasoning, assess alignment with clinical intent, and record findings as formal assurance evidence [11, 13] In doing so, the ECSF reframes explainability as an active safety control, consistent with NHS AI Assurance principles on transparency and human oversight [7, 9].
@@ 4.2 Strengthening Traceability, Validation, and Accountability
By integrating these interpretability outputs into DCB artefacts such as the Hazard Log and Clinical Safety Case, the ECSF embeds traceability and accountability across the lifecycle. This structure reinforces the oversight principles central to the Good Machine Learning Practice (GMLP), the EU Artificial Intelligence Act, and the NHS AI Assurance Framework [9, 10, 7].
@@ 4.3 Addressing Ethical and Human Factors
Explainability must remain clinically usable. Overly complex or inconsistent outputs can increase cognitive burden and erode clinician trust. The ECSF therefore promotes a tiered approach: simplified outputs for clinicians, detailed audit trails for Clinical Safety Officers (CSOs), and technical diagnostics for developers. This layered structure aligns with the NHS AI Assurance Framework’s emphasis on clarity and proportionality, ensuring that interpretability enhances rather than hinders safe clinical adoption [7].
@@ 4.4 Regulatory Convergence and Practical Alignment
For example, the transparency principle of the GMLP is operationalised through ECSF explainability checkpoints, while the EU AI Act’s documentation and oversight obligations are fulfilled through traceable interpretability evidence. Similarly, the NHS AI Assurance Framework’s fairness and accountability requirements are reinforced through bias detection and longitudinal explainability monitoring. By embedding explainability evidence within DCB artefacts such as the Clinical Safety Case, Hazard Log, and Post-Market Surveillance (PMS) Plan, the ECSF ensures that these frameworks function as integrated components of a single safety ecosystem. This convergence reduces duplication, enhances regulatory coherence, and enables local governance teams to meet both national and international assurance expectations simultaneously.
@@ 4.6 Pathways for Implementation
Successful implementation of the Explainability-Enabled Clinical Safety Framework (ECSF) will require coordinated action across three operational domains.
@@
Governance Alignment. Embed explainability checkpoints within DCB0129 and DCB0160 workflows, ensuring they are explicitly referenced in the Hazard Log, Clinical Safety Case, and Post-Market Surveillance (PMS) Plan. Drift thresholds and change-control criteria should be defined in accordance with NHS AI Assurance and Good Machine Learning Practice (GMLP) expectations [7, 9]. These measures ensure that explainability evidence remains aligned with regulatory governance and lifecycle management.
Training and Organisational Culture. Equip Clinical Safety Officers (CSOs), developers, and governance teams with the skills required to interpret and apply explainability outputs in safety management. Fostering a culture of transparency will normalise interpretability as a routine component of clinical safety rather than a technical add-on.
@@ 4.7 Implications for Future Regulation and Research
The ECSF contributes to the emerging international consensus that explainability should be recognised as a regulatory and clinical-safety enabler rather than a purely technical goal. By embedding explainability within the DCB lifecycle, it supports post-market monitoring, human oversight, and traceability requirements articulated in the EU Artificial Intelligence Act and anticipated in future UK Medical Device Regulations [10].
@@ 5 Conclusion
This paper proposes the Explainability-Enabled Clinical Safety Framework (ECSF), which embeds explainability as routine evidence within the DCB0129 and DCB0160 lifecycle. By aligning established interpretability techniques with existing NHS clinical safety artefacts, the framework converts model reasoning into auditable assurance evidence that supports hazard identification, validation and verification, clinical evaluation, risk control, and post-market surveillance. In doing so, it bridges deterministic safety processes with the probabilistic behaviour of AI systems and operationalises the principles of transparency, traceability, and human oversight articulated in current UK and EU guidance [9, 10, 7].
The framework’s practical contribution lies in specifying where and how global and local model-specific explanations populate the Hazard Log, Clinical Safety Case, Clinical Evaluation Report, and Post-Market Surveillance documentation. This strengthens traceability of decision pathways, improves the evidentiary basis of safety arguments, and supports continuous monitoring over time. It also provides a structured mechanism for organisations to meet converging expectations across NHS AI Assurance, Good Machine Learning Practice (GMLP), and the EU AI Act while maintaining full compatibility with current DCB workflows [3, 4, 9, 10].

### v2

@@ 2.1 Framework Development Approach
This study adopted a conceptual modeling approach and cross-regulatory synthesis to integrate explainability principles within existing NHS clinical-safety standards. Rather than proposing a new assurance framework, the analysis sought to align established safety, ethical, and regulatory instruments, including DCB0129 and DCB0160, the GMLP principles, the EU AI Act, and the NHS AI Assurance and T.E.S.T. frameworks [3, 4, 9, 10, 7, 8]. The objective was to identify practical points of convergence where explainability could enhance clinical risk management and governance.
The development of the framework followed three sequential stages. The first stage, regulatory mapping, involved analysing each clause of the DCB0129 and DCB0160 lifecycle in relation to corresponding principles in the GMLP, the EU AI Act, and the NHS AI Assurance frameworks. This mapping process revealed conceptual overlaps such as traceability, validation, and human oversight, as well as areas of divergence including interpretability requirements and post-market transparency obligations. The findings from this exercise provided the structural foundation for the ECSF.
@@
The third stage, framework construction, involved synthesising insights from the regulatory and technical mappings into the framework. The ECSF introduces five explainability checkpoints aligned with the DCB lifecycle: hazard identification, validation and verification, clinical evaluation, risk control, and post-market surveillance. Each checkpoint specifies acceptable forms of explainability evidence, the corresponding DCB artefact, and the assurance purpose such as validation, traceability, or drift monitoring. Collectively, these checkpoints establish explainability as a structured, repeatable, and measurable component of clinical-safety assurance.
@@ 2.2 Analytical Perspectives
To ensure coherence between regulatory intent, technical evidence, and clinical usability, the analysis was organised around three complementary perspectives that together informed the construction of the ECSF.
The first perspective, the regulatory analysis, examined how explainability principles align with external obligations defined by the GMLP guidance, the EU AI Act, and the NHS AI Assurance frameworks [9, 10, 7]. Clauses within DCB0129 and DCB0160 were cross-referenced with these instruments to identify where explainability outputs could provide evidence of compliance with transparency, documentation, and human oversight requirements. Article 13 of the EU AI Act, which mandates transparency and provision of information to deployers, overlaps most strongly with GMLP Principles 1, 9, and 10, and to a lesser extent Principle 3. Together, these principles emphasise multidisciplinary expertise, data representativeness, clear user information, and lifecycle monitoring. This alignment ensures that explainability contributes directly to the evidentiary trail required for regulatory assurance, rather than functioning as a separate technical exercise.
@@ 2.3 Data Sources
The conceptual analysis relied on a combination of regulatory, policy, and academic sources to ensure that the ECSF was grounded in both policy compliance and scientific evidence. The regulatory foundation was established through the review of core NHS and governmental standards, including the NHS Digital clinical-safety standards DCB0129 and DCB0160 [3, 4], the GMLP guiding principles [9], the EU Artificial Intelligence Act [10], and the NHS AI Assurance and T.E.S.T. frameworks published between 2023 and 2024 [7, 8]. These documents provided the structural and procedural requirements against which explainability principles were mapped.
@@ 3.1 Overview
The ECSF integrates established XAI techniques into the DCB0129 and DCB0160 lifecycle to enhance transparency, traceability, and accountability. The framework aligns the deterministic assurance logic of existing NHS safety standards with the probabilistic behaviour of AI systems and reframes explainability from a research objective into a source of verifiable clinical-safety evidence [11, 12, 13].
@@
This approach preserves compliance with existing DCB standards while operationalising the transparency and human-oversight principles articulated in the GMLP guidance and the transparency obligations of the EU AI Act [9, 10].
@@ 3.3 Explainability for Large Language Models and Generative AI
Global LLM explainability. System-level traceability can be achieved through structured documentation and lifecycle transparency. Model-card reporting summarises the model’s intended use, limitations, and performance characteristics, providing a foundation for accountability and auditability [24]. Training-data lineage documentation ensures provenance and representativeness of datasets, aligning with transparency and oversight principles in the GMLP guidance and the EU AI Act [9, 10]. Logging of fine-tuning parameters and model versioning during updates supports reproducibility and traceability throughout iterative development cycles [20]. Finally, specification of guardrail logic within safety documentation maintains human oversight and behavioural control, consistent with NHS AI Assurance and T.E.S.T. frameworks [7, 8].
@@ 3.4 Integration into Assurance Activities and Artefacts
Within the ECSF, explainability evidence is systematically integrated into the five checkpoints of the DCB lifecycle. Each checkpoint aligns with existing assurance artefacts, such as the Hazard Log, Clinical Safety Case, Clinical Evaluation Report, and Post-Market Surveillance documentation, ensuring that transparency, interpretability, and oversight are evidenced throughout the system’s lifecycle. This integration preserves DCB conformance while extending the assurance process to address the probabilistic behaviour of AI systems.
@@
Explainability-Enhanced Validation – Case-Level Explainability (Local). During model validation, local explainability techniques demonstrate whether predictions align with expected clinical reasoning. Methods such as LIME, SHAP, Integrated Gradients, and saliency mapping clarify how specific input variables contribute to individual predictions, providing case-level interpretability that complements quantitative validation metrics [17]. In LLMs, token-level attention visualisation and rationale-trace comparison provide similar evidence. These outputs substantiate the verification and validation activities required under DCB0129 Section 3.5 and are incorporated into the Clinical Safety Case as supporting validation evidence.
Clinical Evaluation – Usability and Human Oversight. Explainability also supports human-factors assurance by enabling clinicians to interrogate model outputs during usability testing. Interactive dashboards, contrastive and counterfactual examples, and rationale summaries allow end users to review AI-generated recommendations, confirming that model reasoning aligns with clinical context and professional judgement [14, 15]. This stage aligns with GMLP Principle 7 on the performance of the human–AI team and Principle 9 on transparency and user information. [9]. Findings are summarised within the Clinical Evaluation Report and, where interpretive or cognitive risks are identified, recorded as updates to the Hazard Log. These activities strengthen the feedback loop between technical validation and clinical usability, reinforcing interpretability as a core dimension of safety.
@@ 3.5 Illustrative Application Scenario
To demonstrate the practical application of the ECSF, a simulated example of an AI early-warning system for sepsis prediction was developed. The model analysed routinely collected vital signs and laboratory data to estimate sepsis risk. This example is illustrative and does not represent a live deployment; rather, it shows how explainability evidence can be systematically embedded within DCB0129 and DCB0160 artefacts across the assurance lifecycle.
Hazard Identification. An agggregate SHAP analysis revealed disproportionate weighting of age and white blood cell count, with under-representation of lactate and respiratory rate, indicating potential model bias and an increased risk of missed atypical presentations. This issue was recorded as a new hazard in the Hazard Log and supported by a FMEA detailing the failure mode, potential clinical impact, and proposed mitigation measures.
@@
Post-Market Surveillance. Within the Post-Market Surveillance (PMS) Plan, periodic audits of explanation stability and feature-importance drift were implemented to detect changes in model reasoning over time, corresponding to ECSF Checkpoint C5 PMS.
@@ 3.6 ECSF Alignment with Regulatory Principles and Key Outcomes
The ECSF aligns with the primary regulatory and ethical instruments governing the safe use of artificial intelligence in the NHS. It operationalises the principles of transparency, traceability, and human oversight established in the GMLP guidelines, the EU AI Act, and the NHS AI Assurance and T.E.S.T. frameworks, while remaining consistent with the DCB0129 and DCB0160 clinical risk management standards [9, 10, 7, 8].
@@ 4.1 Reframing Explainability as a Safety Enabler
The ECSF provides a pragmatic mechanism for translating abstract interpretability research into operational clinical-safety practice. Traditional NHS assurance processes, developed for deterministic software, assume consistent performance under fixed parameters. However, the adaptive and probabilistic behaviour of AI systems challenges these assumptions. Even under controlled conditions, model outputs can vary across runs or datasets due to stochastic optimisation processes and sampling effects [31, 2]. This inherent variability complicates verification and reproducibility, requiring new assurance mechanisms that make uncertainty observable and manageable rather than eliminated.
Explainability provides this mechanism by transforming AI from an opaque computational process into an auditable system. Rather than seeking to remove uncertainty, the ECSF manages it through structured visibility. Explainability checkpoints embedded within the DCB0129 and DCB0160 lifecycle document, interpret, and contextualise model behaviour, converting opacity into evidence. This allows CSOs to trace model reasoning, assess alignment with clinical intent, and record findings as formal assurance evidence [11, 13] In doing so, the ECSF reframes explainability as an active safety control, consistent with NHS AI Assurance principles on transparency and human oversight [7, 9].
@@ 4.2 Strengthening Traceability, Validation, and Accountability
By integrating these interpretability outputs into DCB artefacts such as the Hazard Log and Clinical Safety Case, the ECSF embeds traceability and accountability across the lifecycle. This structure reinforces the oversight principles central to the GMLP, the EU AI Act, and the NHS AI Assurance Framework [9, 10, 7].
@@ 4.3 Addressing Ethical and Human Factors
Explainability must remain clinically usable. Overly complex or inconsistent outputs can increase cognitive burden and erode clinician trust. The ECSF therefore promotes a tiered approach: simplified outputs for clinicians, detailed audit trails for CSOs, and technical diagnostics for developers. This layered structure aligns with the NHS AI Assurance Framework’s emphasis on clarity and proportionality, ensuring that interpretability enhances rather than hinders safe clinical adoption [7].
@@ 4.4 Regulatory Convergence and Practical Alignment
For example, the transparency principle of the GMLP is operationalised through ECSF explainability checkpoints, while the EU AI Act’s documentation and oversight obligations are fulfilled through traceable interpretability evidence. Similarly, the NHS AI Assurance Framework’s fairness and accountability requirements are reinforced through bias detection and longitudinal explainability monitoring. By embedding explainability evidence within DCB artefacts such as the Clinical Safety Case, Hazard Log, and Post-Market Surveillance Plan, the ECSF ensures that these frameworks function as integrated components of a single safety ecosystem. This convergence reduces duplication, enhances regulatory coherence, and enables local governance teams to meet both national and international assurance expectations simultaneously.
@@ 4.6 Pathways for Implementation
Successful implementation of the ECSF will require coordinated action across three operational domains.
@@
Governance Alignment. Embed explainability checkpoints within DCB0129 and DCB0160 workflows, ensuring they are explicitly referenced in the Hazard Log, Clinical Safety Case, and Post-Market Surveillance Plan. Drift thresholds and change-control criteria should be defined in accordance with NHS AI Assurance and GMLP expectations [7, 9]. These measures ensure that explainability evidence remains aligned with regulatory governance and lifecycle management.
Training and Organisational Culture. Equip CSOs, developers, and governance teams with the skills required to interpret and apply explainability outputs in safety management. Fostering a culture of transparency will normalise interpretability as a routine component of clinical safety rather than a technical add-on.
@@ 4.7 Implications for Future Regulation and Research
The ECSF contributes to the emerging international consensus that explainability should be recognised as a regulatory and clinical-safety enabler rather than a purely technical goal. By embedding explainability within the DCB lifecycle, it supports post-market monitoring, human oversight, and traceability requirements articulated in the EU AI Act and anticipated in future UK Medical Device Regulations [10].
@@ 5 Conclusion
This paper proposes the ECSF, which embeds explainability as routine evidence within the DCB0129 and DCB0160 lifecycle. By aligning established interpretability techniques with existing NHS clinical safety artefacts, the framework converts model reasoning into auditable assurance evidence that supports hazard identification, validation and verification, clinical evaluation, risk control, and post-market surveillance. In doing so, it bridges deterministic safety processes with the probabilistic behaviour of AI systems and operationalises the principles of transparency, traceability, and human oversight articulated in current UK and EU guidance [9, 10, 7].
The framework’s practical contribution lies in specifying where and how global and local model-specific explanations populate the Hazard Log, Clinical Safety Case, Clinical Evaluation Report, and Post-Market Surveillance documentation. This strengthens traceability of decision pathways, improves the evidentiary basis of safety arguments, and supports continuous monitoring over time. It also provides a structured mechanism for organisations to meet converging expectations across NHS AI Assurance, GMLP, and the EU AI Act while maintaining full compatibility with current DCB workflows [3, 4, 9, 10].